# ART: Adaptive Resampling-based Training for Imbalanced Classification

### v1

@@ 3.3. Blending with Class Priors
4. Experiments
4.1. Experimental Setup
We evaluate ART against a broad suite of existing methods across five datasets commonly used for imbalanced classification tasks: Pima Indians Diabetes (Smith et al., 1988), Yeast (Nakai, 1991), Red Wine Quality (Cortez et al., 2009), a custom long-tailed version of MNIST (LeCun et al., 1998) (MNIST-LT), and a synthetically imbalanced IMDb sentiment classification dataset (Maas et al., 2011) (IMDb-Custom). MNIST-LT was prepared by taking an imbalance factor of 0.01 and the number of samples for majority class were taken as 1500. All samples were chosen randomly in accordance to the imbalance factor. IMDb-Custom dataset was prepared by transforming the reviews into embedding vectors using all-MiniLM-L6-v2 (Reimers and Gurevych, 2021) by sentence-transformers. Then an arbitrary imbalance ratio of $7$ was chosen and samples were randomly chosen from the data in accordance to it. Pima Indians Diabetes dataset is a binary classification dataset with an imbalance ratio of 1.86. Red Wine Quality dataset is a multi-class classification dataset with 6 classes. It has an imbalance ratios ranging from 1.06 to 68.1 in comparison to the class with most instances. Yeast dataset is a multi-class classification dataset with 9 classes, with imbalance ratios ranging from 1.08 to 23.15 in comparison to the class with highest number of instances. These datasets encompass both binary and multi-class classification scenarios and vary substantially in terms of class imbalance, feature dimensionality, and input modality (tabular, image, and text).
@@ 4.1. Experimental Setup
4.2. Compared Methods
@@
  • No Imbalance Handling (Baseline)
  • ROS
  • RUS
  • SMOTE
@@
  • OHEM
  • LDAM + DRW
4.3. Evaluation Protocol
@@
  • Mean and standard deviation of macro-F1 scores on the held-out test set.
@@
More information about measures taken to ensure fairness and reproducibility in results can be found in Appendix  B
4.4. Test Results
Table 1 presents the macro-F1 scores across all five datasets.
Table 1. Performance comparison of ART and existing methods on Macro-F1 score (mean $\pm$ std over 20 seeds).
@@ 4.4. Test Results
Cost-Sensitive Learning 0.7442 $\pm$ 0.0480 0.3912 $\pm$ 0.0689 0.2790 $\pm$ 0.0266 0.8808 $\pm$ 0.0468 0.722 ± 0.0174
@@
ART achieves the highest macro-F1 scores across all datasets. ART consistently performs well on other metrics, which can be seen in Table  1 and Appendix  C. Below, we provide dataset-specific observations:
Pima Indians Diabetes
ART achieves a consistent improvement over strong baselines, demonstrating its effectiveness in moderately imbalanced binary classification tasks.
Yeast
This dataset poses a significant challenge due to its multi-class nature. ART demonstrates clear superiority, likely due to its adaptive class-wise attention mechanism.
Red Wine Quality
Given its severe class imbalance, this dataset emphasizes ART’s capacity to handle challenging, skewed distributions effectively.
MNIST-LT
ART maintains stable performance, highlighting its robustness in long-tailed image classification scenarios.
IMDb-Custom
ART demonstrates competitive performance, illustrating the flexibility of adaptive resampling techniques in text classification.
To verify that ART’s improvements are not due to random seed variation, we test each baseline against ART with both a paired $t$-test and a two-sided Wilcoxon signed-rank test over the same 20 seeds used in the main experiments. Appendix –A reports the resulting p-values, where entries below 0.05 indicate a statistically significant advantage for ART. We consistently achieve p-value ¡ 0.05 against most methods, indicating significant performance gains when using ART compared to other methods.
4.5. Ablation Studies
@@
4.5.1. Effect of the Blending Constant $c$
@@
Figure 3. Comparison of ART performance over different blending constant ($c$) where c is a hyperparameter. Results show ART is insensitive to the choice of $c$ and requires minimal hyperparameter optimization.
[…]
The image shows a line plot comparing the performance of ART on macro F1-score with different values of c over different bf values.
4.5.2. Effect of Boost Frequency bfbf
@@
Figure 4. Comparison of performance of ART across varying boosting frequency (bfbf) values while keeping all other hyperparameters constant. Experiments over varying c values show performance is insensitive to choice of boosting frequency.
[…]
The image shows line plot comparing the performance of changing boosting frequency in ART. It shows 3 lines which show change in performance over blending constant values 0.25, 0.5 and 0.75 respectively. For each line, boosting frequency is evaluated to a range of 1 to 8. Performance is compared over macro F1-score.
4.5.3. Effect of Model Width
@@
Figure 5. Comparison of baseline and ART performance over different model width. Other model architecture choices and training hyperparameters were kept constant.
[…]
The graph shows a line plot comparing the performance on macro F1-score by changing the model width. It has 2 lines showing the performance changes of baseline method and ART respectively.
4.5.4. Performance Across Imbalance Ratios
@@
As shown in Figure 6, ART consistently outperforms the baseline across all imbalance ratios. While both methods experience a decline in macro-F1 as imbalance severity increases, ART demonstrates significantly better stability. For instance, at an imbalance ratio of 10, ART achieves a macro-F1 of approximately 0.68, whereas the baseline drops to around 0.56. Even at extreme imbalance levels, ART maintains a notable advantage, highlighting its ability to adapt sampling in response to increasing class difficulty.
Figure 6. ART maintains higher performance than Baseline at most imbalance levels, with both methods showing a decline as imbalance increases. Shaded areas indicate variability across runs.
[…]
The graph shows a line plot comparing the performance on macro F1-score by changing the imbalance ratios. It has 2 lines showing the performance changes of baseline method and ART respectively.
5. Conclusion
@@
5.1. Results
Across five benchmark datasets covering tabular, image, and text modalities, ART consistently achieved the strongest macro-F1 scores, outperforming every method we tested. On the Pima, Yeast, and Red Wine datasets, it improved over the best competitor by 1.0–3.8 percentage points and increased the mean macro-F11 across all tabular datasets by 2.64 percentage points. These gains were statistically significant according to paired $t$-tests and Wilcoxon signed-rank tests ($p<0.05$). ART also maintained its advantage on the long-tailed MNIST-LT and the synthetically imbalanced IMDb-Custom dataset, confirming its modality-agnostic design. Notably, ART matched or exceeded baseline performance even when the underlying model’s width was reduced by a factor of four, confirming that it can maintain high performance under limited computational resources.
5.2. Limitations
ART assumes an epoch-based training loop that (i) computes batch-wise losses, (ii) aggregates gradients, and (iii) evaluates class-wise metrics on a validation set at regular intervals. Thus, it applies well to models trained with stochastic gradient descent (e.g. MLPs, and CNNs) but is unsuitable for non-iterative learners, like decision trees, or k-NN. In addition, computing class-wise F1 every bf epochs incurs overhead proportional to the validation-set size; for very large datasets, this cost may be non-trivial unless subsampling is employed.
5.3. Future Work
@@ Appendix A Statistical Significance Tests
Table 2. Pima Dataset: p-values (ART vs. baselines)
@@
Table 3. Yeast Dataset: p-values (ART vs. baselines)
@@
Table 4. Red Wine Dataset: p-values (ART vs. baselines)
@@
Table 5. MNIST-LT Dataset: p-values (ART vs. baselines)
@@
Table 6. IMDb-Custom Dataset: p-values (ART vs. baselines)
@@ Appendix B Reproducibility Details
To facilitate exact replication of our results, we provide below the fixed random-seed list, hardware configuration, and software environment used for every experiment in Section 4.
B.1. Random Seeds
@@ Appendix C Performance Metrics
Table 7. Performance comparison of ART and existing methods on Accuracy (mean $\pm$ std over 20 seeds).
@@
Table 8. Performance comparison of ART and existing methods on Precision (mean $\pm$ std over 20 seeds).
@@
Table 9. Performance comparison of ART and existing methods on Recall (mean $\pm$ std over 20 seeds).

### v2

@@ 3.4 Blending with Class Priors
3.5 Computational Overhead
ART differs from the baseline only in the periodic refresh step that updates the sampling distribution and rebuilds the training loader. The baseline trains on a fixed training set and does not perform any refresh. Let $E$ denote the number of epochs actually executed and let $bf$ be the boost frequency. ART triggers a refresh every $bf$ epochs, so the number of refresh events is
$R=\Big\lfloor\frac{E}{bf}\Big\rfloor.$
We express the total runtime as
$T_{ART}=T_{baseline}+\Big\lfloor\frac{E}{bf}\Big\rfloor\cdot T_{refresh},$
where $T_{baseline}$ is the baseline wall-clock time for training and validation over $E$ epochs, and $T_{refresh}$ is the wall-clock time of one refresh event.
A refresh event consists of three components:
$T_{refresh}=T_{val\_fwd}+T_{metric}+T_{resample}.$
Here,
  • 1.
    $T_{val\_fwd}$ is the time for a single forward pass over the validation set.
  • 2.
    $T_{metric}$ is the time to compute per-class F1 scores from the validation predictions and calculate the class weights.
  • 3.
    $T_{resample}$ is the time to select the indices for resampling based on the class weights.
Space overhead
ART maintains the same asymptotic dataset size as the baseline. It allocates new arrays when rebuilding the refreshed dataset, so peak memory usage may increase during a refresh, but no state proportional to the number of epochs is stored.
4 Experiments
4.1 Experimental Setup
We evaluate ART against a broad suite of existing methods across five datasets commonly used for imbalanced classification tasks: Pima Indians Diabetes (Smith et al., 1988), Yeast (Nakai, 1991), Red Wine Quality (Cortez et al., 2009), a custom long-tailed version of MNIST (LeCun et al., 1998) (MNIST-LT), and a synthetically imbalanced IMDb sentiment classification dataset (Maas et al., 2011) (IMDb-Custom). MNIST-LT was prepared with an imbalance factor of 0.01, and the number of samples for the majority class was set to 1500. All samples were randomly selected according to the imbalance factor. IMDb-Custom dataset was prepared by transforming the reviews into embedding vectors using all-MiniLM-L6-v2 (Reimers and Gurevych, 2021) by sentence-transformers. Then an arbitrary imbalance ratio of $7$ was chosen and samples were randomly chosen from the data in accordance to it. Pima Indians Diabetes dataset is a binary classification dataset with an imbalance ratio of 1.86. Red Wine Quality dataset is a multi-class classification dataset with 6 classes. It has an imbalance ratio ranging from 1.06 to 68.1 in comparison to the class with the most instances. The yeast dataset is a multi-class classification dataset with 9 classes, with imbalance ratios ranging from 1.08 to 23.15 in comparison to the class with the highest number of instances. These datasets encompass both binary and multi-class classification scenarios and vary substantially in terms of class imbalance, feature dimensionality, and input modality (tabular, image, and text).
@@ 4.1 Experimental Setup
4.1.1 Macro F1 as an Appropriate Control Signal for ART
Choosing the right evaluation metric for adaptive resampling is important because it affects how training resources are distributed among classes. In ART, we use class-wise macro F1 to adjust the sampling distribution for both practical and methodological reasons.
First, macro F1 measures the balance between precision and recall for each class. In imbalanced classification, minority classes often have more false negatives, which lowers recall but does not affect overall accuracy much. Unlike accuracy or cross-entropy loss, class-wise F1 directly penalizes these errors, making it better for improving the recognition of minority classes. Because macro F1 treats all classes equally, it helps identify underperforming classes regardless of how often they appear.
Second, using macro F1 as feedback aligns with the adaptive sampling strategy and the evaluation metric commonly used in imbalanced learning. Many real-world tasks, such as medical diagnosis, fraud detection, and rare-event prediction, prioritize balanced performance over overall accuracy. By using class-wise F1, ART ensures the resampling process uses the same metric used to evaluate model performance, reducing any mismatch between training and evaluation.
Third, using metric-based feedback is more stable than relying on instance-level difficulty. Signals based on loss values or margins can be affected by outliers, noisy labels, and random batch effects. In contrast, class-wise F1 is calculated over the whole validation set, giving a more reliable estimate of class difficulty. This approach helps avoid overreacting to single hard or mislabeled examples, which is a common problem with instance-level sampling.
Finally, macro F1 is easy to use and works with any model. It can be calculated for any classifier and does not depend on the loss function or model structure. This means ART can be used with different types of models and data, as shown in our experiments with tabular, image, and text datasets.
In summary, macro F1 provides a stable, evaluation-focused signal for adaptive resampling. By using this metric at the class level, ART can focus training on classes that need improvement and avoid the instability and complexity of instance-level difficulty modeling.
4.2 Compared Methods
@@
  • 1.
    Random Oversampling (ROS)
  • 2.
    Random Undersampling (RUS)
  • 3.
    Synthetic Minority Over-sampling Technique (SMOTE)
  • 4.
@@
  • 3.
    Online Hard Example Mining (OHEM)
  • 4.
    Label-Distribution-Aware Margin + Deferred Re-Weighting (LDAM + DRW)
Hybrid Methods:
  • 1.
    Balanced Meta-Softmax (BALMS)
4.3 Evaluation Protocol
@@
  • 1.
    Mean and standard deviation of macro F1 scores on the held-out test set.
  • 2.
@@
More information about measures taken to ensure fairness and reproducibility in results can be found in B
4.4 Test Results
Table 1 presents the macro F1 scores across all five datasets.
Table 1: Performance comparison of ART and existing methods on macro F1 score (mean $\pm$ std over 20 seeds).
@@ 4.4 Test Results
Loss-based Methods
Cost-Sensitive Learning 0.7442 $\pm$ 0.0480 0.3912 $\pm$ 0.0689 0.2790 $\pm$ 0.0266 0.8808 $\pm$ 0.0468 0.722 $\pm$ 0.0174
@@
ART achieves the highest macro F1 scores across all datasets, which is visualized in Figure 3 using a Delta matrix heatmap that compares ART and competing methods across all datasets.
Figure 3: Delta-Matrix heatmap showing the difference in macro F1 scores between ART and competing methods across all datasets. Positive (green) values indicate improvements of ART over the compared method, while negative (red) values denote cases where the competing method performs better.
ART consistently performs well on other metrics, as shown in Table  1 and C. In the Pima Indians Diabetes dataset, ART consistently outperforms strong baselines, demonstrating its effectiveness in moderately imbalanced binary classification tasks. The Yeast dataset poses a significant challenge due to its multi-class nature. ART demonstrates clear superiority, likely due to its adaptive class-wise attention mechanism. The Red Wine quality dataset, with its severe class imbalance, emphasizes ART’s capacity to handle challenging, skewed distributions effectively. ART maintains stable performance, highlighting its robustness in long-tailed image classification scenarios like the MNIST-LT dataset. It demonstrates competitive performance, illustrating the flexibility of adaptive resampling techniques in text classification of the IMDb-custom dataset.
To verify that ART’s improvements are not due to random seed variation, we test each baseline against ART with both a paired $t$-test and a two-sided Wilcoxon signed-rank test over the same 20 seeds used in the main experiments. A reports the resulting p-values, where entries below 0.05 indicate a statistically significant advantage for ART. We consistently achieve p-value ¡ 0.05 against most methods, indicating significant performance gains when using ART compared to other methods.
4.5 Ablation Studies
@@
4.5.1 Effect of the Blending Constant $c$
@@
Figure 4: Comparison of ART performance over different blending constant ($c$) where c is a hyperparameter. Results show ART is insensitive to the choice of $c$ and requires minimal hyperparameter optimization.
4.5.2 Effect of Boost Frequency bfbf
@@
Figure 5: Comparison of performance of ART across varying boosting frequency (bfbf) values while keeping all other hyperparameters constant. Experiments over varying c values show performance is insensitive to choice of boosting frequency.
4.5.3 Effect of Model Width
@@
Figure 6: Comparison of baseline and ART performance over different model width. Other model architecture choices and training hyperparameters were kept constant.
4.5.4 Performance Across Imbalance Ratios
@@
As shown in Figure 7, ART consistently outperforms the baseline across all imbalance ratios. While both methods experience a decline in macro F1 as imbalance severity increases, ART demonstrates significantly better stability. For instance, at an imbalance ratio of 10, ART achieves a macro F1 of approximately 0.68, whereas the baseline drops to around 0.56. Even at extreme imbalance levels, ART maintains a notable advantage, highlighting its ability to adapt sampling in response to increasing class difficulty.
Figure 7: ART maintains higher performance than Baseline at most imbalance levels, with both methods showing a decline as imbalance increases. Shaded areas indicate variability across runs.
5 Conclusion
@@
5.1 Results
Across five benchmark datasets covering tabular, image, and text modalities, ART consistently achieved the strongest macro F1 scores, outperforming every method we tested. On the Pima, Yeast, and Red Wine datasets, it improved over the best competitor by 1.0–3.8 percentage points and increased the mean macro-F11 across all tabular datasets by 2.64 percentage points. These gains were statistically significant according to paired $t$-tests and Wilcoxon signed-rank tests ($p<0.05$). ART also maintained its advantage on the long-tailed MNIST-LT and the synthetically imbalanced IMDb-Custom dataset, confirming its modality-agnostic design. Notably, ART matched or exceeded baseline performance even when the underlying model’s width was reduced by a factor of four, confirming that it can maintain high performance under limited computational resources.
5.2 Limitations
ART assumes an epoch-based training loop that (i) computes batch-wise losses, (ii) aggregates gradients, and (iii) evaluates class-wise metrics on a validation set at regular intervals. Thus, it applies well to models trained with stochastic gradient descent (e.g. MLPs, and CNNs) but is unsuitable for non-iterative learners, like decision trees, or k-NN. In addition, computing class-wise F1 every bf epochs incurs overhead proportional to the validation-set size; for very large datasets, this cost may be non-trivial unless subsampling is employed. ART relies on a held-out validation set to estimate class-wise performance, which may reduce effective training data in extremely small datasets. Additionally, because low F1 scores may arise from label noise rather than intrinsic difficulty, ART may overemphasize noisy classes; mitigating this effect is a direction for future work.
5.3 Future Work
@@
Declarations:
The code is available online at Github Link.
@@ Appendix A Statistical Significance Tests
Table 2: Pima Dataset: p-values (ART vs. baselines)
@@
BALMS 0.0010 0.0020
Table 3: Yeast Dataset: p-values (ART vs. baselines)
@@
BALMS 0.0133 0.0136
Table 4: Red Wine Dataset: p-values (ART vs. baselines)
@@
BALMS 0.0717 0.1893
Table 5: MNIST-LT Dataset: p-values (ART vs. baselines)
@@
BALMS 0.8970 0.9854
Table 6: IMDb-Custom Dataset: p-values (ART vs. baselines)
@@ Appendix B Reproducibility Details
To facilitate exact replication of our results, we provide below the fixed random-seed list, hardware configuration, and software environment used for every experiment in Section 4. The code is available online at Github Link.
B.1 Random Seeds
@@ Appendix C Performance Metrics
Table 7: Performance comparison of ART and existing methods on Accuracy (mean $\pm$ std over 20 seeds).
@@
Table 8: Performance comparison of ART and existing methods on Precision (mean $\pm$ std over 20 seeds).
@@
Table 9: Performance comparison of ART and existing methods on Recall (mean $\pm$ std over 20 seeds).